OpenStudy (anonymous):

which expression is equivalent to (x+3)(3x^2-4x-5)

OpenStudy (anonymous):

3x^3+5x^2-17x-15 x(x(3x+5)-17)-15 (x+3)(x(3x-4)-5) x=-3 x=1/3)(2- Sqroot 19 x=1/3(2+ Sroot 19

OpenStudy (anonymous):

Thanks!